OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12 Answer x = –2 and x = 10 x = 2 and x = –10 x = –2 and x = –10 x = 2 and x = 10

OpenStudy (anonymous):

x=-2 and x=10

OpenStudy (ksaimouli):

1

OpenStudy (anonymous):

|–2x + 8| = 12 –2x + 8 = 12 8 - 12 = 2x -> x = -4/2 = -2 –2x + 8 = -12 8 + 12 = 2x -> x = 20/2 = 10 Thus x = -2, x =10